Solve the following system of equations, using the elimination method:
x - 2y = -12
2x + 9y = 2

Answer by TheInstructor(29)

Last question for tonight!
Okay elimination method involves tweaking the equation so that it suits your purpose. Now it's completely upto you which variable you wish to find first, however it will be most obvious (and easier) which one to solve for first using the elimination method!
x-2y = -12
2x+9y = 2
Multiply the first equation by 2
2(x-2y=-12) = 2x-4y=-24
Now:
2x - 4y = -24
-
2x + 9y = 2
Now you can see that this allows us to get rid of the x variable so:
-4y - 9y = -24 -2
-13y = -26
Therefore y =2
Sub y= 2 into any of your equations:
2x + 9(2) = 2
2x = 2-18
2x = -16
Therefore x = -8
Check it by subbing it into the other equation:
(-8) - 2(2) = -12
Therefore these variables are correct! Hope this made sense.
Stay awesome bro!